The following table shows the actual demand observed over the last 11 years Year 10 11 Demand 7 4 Using exponential smoothing with α = 0.30 and a forecast for year 1 of 6.0, provide the forecast from periods 2 through 12 round your responses place) one decimal Year 2 4 10 12 Forecast 6.0
Answer #1

Calculation of Forecast Using exponential smoothing for the period 2 to 12

Under exponential smoothing, the forecast for next period =

Forecast for previous period (Ft-1)+ Alpha (α) x {Actual of previous period (At-1) – Forecast of previous period(Ft-1)}

Ft = Ft-1 + α(At-1 - Ft-1)

All Forecast Values Round to 1 decimal

Year Demand (A) Forecast (F) Calculations (Only for Explanation
1 7 6
2 9 6.3 6 + .30(7- 6) = 6.30
3 4 7.1 6.3 + .30(9 - 6.3) = 7.11
4 7 6.2 7.1 + .30(4 - 7.1) = 6.17
5 13 6.4 6.2 + .30(7 - 6.2) = 6.44
6 7 8.4 6.4 + .30(13 - 6.4) = 8.38
7 11 8.0 8.4 + .30(7 - 8.4) = 7.98
8 14 8.9 8.0 + .30(11 - 8.0) = 8.9
9 9 10.4 8.9 + .30(14 - 8.9) = 10.43
10 11 10.0 10.4 + .30(9 - 10.4) = 9.98
11 7 10.3 10.0 + .30(11-10) = 10.30
12 9.3 10.3 + .30(7-10.30) = 9.31